# Low-Rank Bandits via Tight Two-to-Infinity Singular Subspace Recovery

### v1

@@ Appendix B Singular Subspace Recovery
In this appendix, we give the proofs of our results on the singular subspace recovery in the two-to-infinity norm. In Appendix B.1, we provide the proof of Theorem 4.1 which in fact follows from Theorem B.1. The latter theorem is also a guarantee on the two-to-infinity norm but where the estimation error is evaluated in an alternative way. The proof of Theorem B.1 is provided in Section B.2 and features two key ideas, namely the leave-one-out analysis (Abbe et al., 2020; Chen et al., 2021) and the Poisson approximation argument (Stojanovic et al., 2023).
@@
First, we present two key results from which Theorem 4.1 follows. We defer proofs of this two claims to Appendices B.2 and B.3, respectively.
@@
$\max\left(\|V-\widehat{V}(\widehat{V}^{\top}V)\|_{2\to\infty},\|U-\widehat{U}(\widehat{U}^{\top}U)\|_{2\to\infty}\right)\leq C\mu^{3}\kappa^{2}r^{3/2}\sqrt{\frac{m+n}{T\omega_{\min}mn(m\wedge n)}\log^{3}\left(\frac{(m+n)T}{\delta}\right)}$
@@
$T\geq\frac{c\mu^{4}\kappa^{2}r^{2}(m+n)}{\omega_{\min}mn}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$
@@
Under the event that $\|\widetilde{M}-M\|_{\textup{op}}\leq\sigma_{r}(M)/2$, we have that
@@
Now the proof of Theorem 4.1 follows straightforwardly. First, we note that under the event that $\|\widetilde{M}-M\|_{\textup{op}}\leq\sigma_{r}(M)/2$, by Lemma B.2 we need to upper bound only the terms $\|U-\widehat{U}\widehat{U}^{\top}U\|_{2\to\infty},\|V-\widehat{V}\widehat{V}^{\top}V\|_{2\to\infty},\|\widetilde{M}-M\|_{\textup{op}}$. Now, observe that we may use Theorem B.1 to upper bound the terms $\|U-\widehat{U}\widehat{U}^{\top}U\|_{2\to\infty}$ and $\|V-\widehat{V}\widehat{V}^{\top}V\|_{2\to\infty}$ with high probability, and use Proposition B.5 to upper bound the term $\|\widetilde{M}-M\|_{\textup{op}}$ with high probability. The obtained upper bounds combined with the fact that $\|M\|_{\max}/\sigma_{r}(M)\leq\mu^{2}\kappa r/\sqrt{mn}$ from Lemma B.8, and $\|U\|_{2\to\infty}\vee\|V\|_{2\to\infty}\leq\mu\sqrt{r/(m\wedge n)}$ yield the final result.
@@
The proof of Theorem B.1 is rather involved but follows the same steps as in the proof of Lemma 30 in (Stojanovic et al., 2023), with a slight difference that in (Stojanovic et al., 2023) authors assume that $\omega_{i,j}=1/(mn)$ for all $i,j$. However, for the sake of completeness we highlight the key steps of the proof next.
@@
For the sake of analysis, it is rather convenient to describe the estimate $\widetilde{M}$ given in (1) with an alternative random matrix which has the same distribution. For all $(i,j)\in[m]\times[n]$, define $Z_{i,j}=\sum_{t=1}^{T}\mathds{1}_{\{(i_{t},j_{t})=(i,j)\}}$ and let $(\xi_{i,j,t}^{\prime})$ be a sequence of i.i.d. random variables that have the same distribution as $\xi_{1}$. Now, observe that our matrix $\widetilde{M}$ has the same distribution as the random matrix $\widetilde{M}^{\prime}$ defined as:
@@
Let $(\Omega,\mathcal{F},\mathbb{P})$ (resp. (Ω,,)normal-$(\Omega,\mathcal{F},\mathbb{P}^{\prime})$) be the probability space under the random matrix model (17) (resp. (18)). Then for any event $\mathcal{E}\in\mathcal{F}$, we have
@@
The proof of Lemma B.3 follows from Lemma B.4 and is similar to the proof of Lemma 20 in (Stojanovic et al., 2023).
@@
We borrow the following Lemma from (Stojanovic et al., 2023) (see their Lemma 14) which itself is a mild generalization of Theorem 5.7 in (Mitzenmacher & Upfal, 2017)
@@
Let $Y_{i}^{(t)}\sim\mathrm{Poisson}(tp_{i})$, $i=1,\dots,n$, be independent random variables with $\sum_{i=1}^{n}p_{i}=1$. Moreover, let $(Z_{1}^{(t)},Z_{2}^{(t)},\dots,Z_{n}^{(t)})\sim\allowbreak\mathrm{Multinomial}(t,(p_{1},\dots,p_{n}))$. Let $f:\mathbb{R}^{p}\to\mathbb{R}_{+}$ be any non-negative function. Then:
@@
The following propositions are direct consequences of a truncated matrix Bernstein inequality (see e.g., Theorem A.3 by (Hopkins et al., 2016)).
@@
Let $A$ be a $m\times 2r$ deterministic matrix, and $B$ be a $n\times 2r$ deterministic matrix. Then, under the random matrix model (18) with compound Poisson entries, and denoting $L=\max(\|M\|_{\max},1)$, we have:
@@
    for all $\ell\in[m]$, for all $\delta\in(0,1)$, for all $T\geq(1/(n\omega_{\min}))\log^{3}(en/\delta)$, the following event
@@
    for all $k\in[n]$, for all $\delta\in(0,1)$, for all $T\geq(1/(m\omega_{\min}))\log^{3}(em/\delta)$, the following event
@@
The proofs of Proposition B.5 and Proposition B.6 are essentially the same as those of Proposition 26 and Proposition 27 in (Stojanovic et al., 2023), and are omitted. Note that both Propositions in (Stojanovic et al., 2023) were derived under i.i.d. sub-gaussian noise assumption.
@@
This part of the analysis follows the steps of the proof of Lemma 30 in (Stojanovic et al., 2023). Therefore, we will only highlight the key ideas and differences.
@@
We can define $\widetilde{S}$ in a similar way, and denote $\widehat{Q}\in\mathbb{R}^{(m+n)\times 2r}$, the matrix of containing the $2r$ eigenvectors of the best $2r$-rank approximation of $\widetilde{S}$. Here, we observe that:
@@
(ii) Second, it can be shown, under the event $\mathcal{E}_{1}=\{\|E\|_{\textup{op}}\leq C_{1}\sigma_{r}(M)\}$ where $C_{1}$ is a universal and sufficiently small positive constant, that:
$\|Q-\widehat{Q}(\widehat{Q}^{\top}Q)\|_{2\to\infty}\leq\frac{1}{\sigma_{2r}(S)}\bigg{(}\frac{5\|SQ\|_{2\to\infty}\|E\|}{\sigma_{2r}(S)}+3\|EQ\|_{2\to\infty}+2\|E(Q-\widehat{Q}(\widehat{Q}^{\top}Q)\|_{2\to\infty}\bigg{)}.$ (22)
@@
Now, the key observation that underpins the leave-one-out analysis is that $E_{\ell,:}$ and $Q-\widehat{Q}^{(\ell)}W_{\widehat{Q}^{(\ell)}}$ are independent provided the entries of the error matrix $E$ are independent. This will be ensured thanks to the Poisson approximation (using Lemma B.3) by controlling the quantities of interest under the random matrix model (18). Define
$\mathcal{B}(\delta)=\sqrt{\frac{(m+n)\log^{3}((m+n)/\delta)}{T\omega_{\min}}}$
@@
$\mathcal{E}_{2}$ =$=\left\{\|E\|_{\textup{op}}\leq C_{2}(\|M\|_{\max}+1)\mathcal{B}(\delta)\right\}$
$\forall\ell\in[m+n],\qquad\mathcal{E}_{3}^{(\ell)}$ =$=\left\{\|E_{\ell,:}(Q-\widehat{Q}^{(\ell)}W_{\widehat{Q}^{(\ell)}})\|_{2}\leq C_{3}(\|M\|_{\max}+1)\|Q-\widehat{Q}^{(\ell)}W_{\widehat{Q}^{(\ell)}}\|_{2\to\infty}\mathcal{B}(\delta)\right\}$
@@
for $c>0$ a universal constant sufficiently large. Additionally, using the fact that $\|M\|_{\max}/\sigma_{r}(M)\leq\mu^{2}\kappa r/\sqrt{mn}$ from Lemma B.8, we can deduce that $\mathcal{E}_{1}\supseteq\mathcal{E}_{2}$ and $\mathbb{P}(\mathcal{E}_{2})\geq 1-\delta$ as well as $\mathbb{P}(\mathcal{E}_{3}^{(\ell)})\geq 1-\delta$ for all $\ell\in[m+n]$ when
$T\geq\frac{c^{\prime}\mu^{4}\kappa^{2}r^{2}(m+n)}{\omega_{\min}mn}\log^{3}\left(\frac{m+n}{\delta}\right)$ (25)
for a universal constant $c^{\prime}$ sufficiently large. At this point, what remains to be done is to deal with the inequalities (22), (23) and (24) under the intersection of the events $\mathcal{E}_{2},\mathcal{E}_{3}^{(1)},\dots,\mathcal{E}_{3}^{(m+n)}$. We spare the readers the details of this tedious task and write directly the resulting bound
$\|Q-\widehat{Q}(\widehat{Q}^{\top}Q)\|_{2\to\infty}\leq C^{\prime}\mu^{3}\kappa^{2}r^{3/2}\frac{\mathcal{B}(\delta)}{\sqrt{mn(m\wedge n)}}$
for some universal constant $C^{\prime}>0$. Thus, by union bound, the above bound holds with probability at least $1-(1+m+n)\delta$, provided condition (25) holds, under the random matrix model (18). Using Lemma B.3, we conclude that the same event holds with probability $1-(1+m+n)e\sqrt{T}\delta$, provided condition (25) holds, under the random matrix model (17). Re-parametrizing by $\delta^{\prime}=e(m+n+1)\sqrt{T}\delta$ concludes the proof.
@@ B.4 Subspace recovery in Frobenius norm
Deriving such guarantees is by now well understood under general noise assumptions (see e.g., (Keshavan et al., 2009; Chen et al., 2021)). Indeed, the classical proof follows by first deriving upper bounds on $\|\widetilde{M}-M\|_{\textup{op}}$ with high probability, then using Davis-Kahan’s or Wedin’s theorem to obtain upper bounds on the subspace recovery error in the Frobenius norm. The following lemma provides upper bounds of the recovery error in the Frobenius norm when sampling under $\omega$ is the uniform distribution over $[m]\times[n]$.
@@
Let $\delta\in(0,1)$. Assume that $p$ is the uniform distribution over $[m]\times[n]$. Then, with probability at least $1-\delta$ holds that
$\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{\textup{F}},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{\textup{F}}\right)\lesssim\mu^{2}\kappa r^{3/2}\sqrt{\frac{m+n}{T}\log^{3}\left(\frac{(m+n)T}{\delta}\right)},$
provided that $T\gtrsim(m+n)\log^{3}\left(\frac{(m+n)T}{\delta}\right)$.
@@
An analogous bound holds for $\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{F}$. Lastly using Proposition B.5 to bound $\|M-\widetilde{M}\|_{\textup{op}}$ and Lemma B.8 to bound $\|M\|_{\max}/\sigma_{r}(M)$ concludes the proof. ∎
@@ B.5 Additional Lemmas
The following result is immediate (see Lemma 17 in (Stojanovic et al., 2023)).
@@ Appendix C Policy Evaluation for Contextual Linear Bandits
In this section, we study the simpler contextual linear bandit problem, whose error bound analysis will be useful to handle the contextual low-rank bandit problem. We also demonstrate that the linear bandit error bound is instance-optimal by deriving a sample complexity lower bound of independent interest.
We observe $T$ i.i.d. samples of the form $(i_{t},j_{t},r_{t})_{1\leq t\leq T}$ where the context $i_{t}\in[m]$ is drawn independently from a known distribution $\rho$ and the arm $j_{t}\in[n]$ is sampled from $\pi^{b}(.|i_{t})$ for a known behavioral policy $\pi^{b}$. The reward distribution is assumed to have a linear structure with respect to a known feature map $\psi:[m]\times[n]\to\mathbb{R}^{d}$ : there exists an unknown parameter $\theta$ in $\mathbb{R}^{d}$ such that $r_{t}=\psi_{i_{t},j_{t}}^{\top}\theta+\xi_{t}$ where $\left(\xi_{t}\right)_{t\leq 1}$ is an i.i.d. sequence of 1-subgaussian noise, independent from all the other random variables. The overall bandit model described above will be denoted as $\mathcal{M}_{\theta}$. For a given target policy $\pi$, we use the previously defined notations $\omega_{i,j}=\rho(x)\pi^{b}(j|i)$, $\omega^{\pi}_{i,j}=\rho(x)\pi(j|i)$ and we define
@@
For any $\delta\in(0,1)$ and any $\varepsilon>0$, we say that $\hat{v}^{\pi}$ [^10] We will often omit the superscript $\pi$ to ease the notations. is $(\varepsilon,\delta)$-PAC if $\mathbb{P}_{\theta}\left(|v^{\pi}_{\theta}-\hat{v}^{\pi}|\leq\varepsilon\right)\geq 1-\delta$ for every $\theta\in\mathbb{R}^{d}$. We will write $v^{\pi}$ for $v^{\pi}_{\theta}$ when there is no ambiguity about the underlying parameter. We also write $N\succcurlyeq M$ (resp. $N\preccurlyeq M$) when $N-M$ (resp. $M-N$) is positive semi-definite, and denote the Moore-Penrose pseudoinverse (Barata & Hussein, 2011) of $N$ by $N^{\dagger}$.

### v2

@@ Appendix B Singular Subspace Recovery
In this appendix, we give the proofs of our results on the singular subspace recovery in the two-to-infinity norm. In Appendix B.1, we provide the proof of Theorem 4.1 which in fact follows from Theorem B.1. The latter theorem is also a guarantee on the two-to-infinity norm but where the estimation error is evaluated in an alternative way. The proof of Theorem B.1 is provided in Appendix B.2 and features two key ideas, namely the leave-one-out analysis (Abbe et al., 2020; Chen et al., 2021) and the Poisson approximation argument (Stojanovic et al., 2024).
@@
We now present the two key results from which Theorem 4.1 follows. We defer the proofs of these two claims to Appendices B.2 and B.3, respectively.
@@
$\max\left(\|U-\widehat{U}(\widehat{U}^{\top}U)\|_{2\to\infty},\|V-\widehat{V}(\widehat{V}^{\top}V)\|_{2\to\infty}\right)\leq C\frac{L}{\sigma_{r}(M)}\sqrt{\frac{\mu^{2}\kappa^{2}r(m+n)}{T\omega_{\min}(m\wedge n)}\log^{3}\left(\frac{(m+n)T}{\delta}\right)},$
@@
$\quad T\geq c\frac{L^{2}(m+n)}{\sigma_{r}^{2}(M)\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$
@@
Under the event $\|\widetilde{M}-M\|_{\textup{op}}\leq\sigma_{r}(M)/2$, we have
@@
Now the proof of Theorem 4.1 follows straightforwardly. First, we note that under the event that $\|\widetilde{M}-M\|_{\textup{op}}\leq\sigma_{r}(M)/2$, by Lemma B.2 we only need to upper bound the terms $\|U-\widehat{U}\widehat{U}^{\top}U\|_{2\to\infty},\|V-\widehat{V}\widehat{V}^{\top}V\|_{2\to\infty},\|\widetilde{M}-M\|_{\textup{op}}$. Now, observe that we may use Theorem B.1 to upper bound $\|U-\widehat{U}\widehat{U}^{\top}U\|_{2\to\infty}$ and $\|V-\widehat{V}\widehat{V}^{\top}V\|_{2\to\infty}$ with high probability, and use Proposition B.5 to upper bound the term $\|\widetilde{M}-M\|_{\textup{op}}$ with high probability. The obtained upper bounds combined with the fact that $\|U\|_{2\to\infty}\vee\|V\|_{2\to\infty}\leq\mu\sqrt{r/(m\wedge n)}$ yield the final result.
@@
The proof of Theorem B.1 is rather involved but follows the same steps as the proof of Lemma 30 in (Stojanovic et al., 2024), with the slight difference that in (Stojanovic et al., 2024), it is assumed that $\omega_{i,j}=1/(mn)$ for all $i,j$. For the sake of completeness, we highlight the key steps of the proof next.
@@
For the sake of the analysis, it is rather convenient to describe the estimate $\widetilde{M}$ given in (1) with an alternative random matrix which has the same distribution. For all $(i,j)\in[m]\times[n]$, define $Z_{i,j}=\sum_{t=1}^{T}\mathds{1}_{\{(i_{t},j_{t})=(i,j)\}}$ and let $(\xi_{i,j,t}^{\prime})$ be a sequence of i.i.d. random variables that have the same distribution as $\xi_{1}$. Now, observe that our matrix $\widetilde{M}$ has the same distribution as the random matrix $\widetilde{M}^{\prime}$ defined as:
@@
Let $(\Omega,\mathcal{F},\mathbb{P})$ (resp. $(\Omega,\mathcal{F},\mathbb{P}^{\prime})$) be the probability space under the random matrix model (13) (resp. (14)). Then for any event $\mathcal{E}\in\mathcal{F}$, we have
@@
The proof of Lemma B.3 follows from Lemma B.4 and is similar to the proof of Lemma 20 in (Stojanovic et al., 2024).
@@
We borrow the following Lemma from (Stojanovic et al., 2024) (see their Lemma 14) which itself is a mild generalization of Theorem 5.7 in (Mitzenmacher & Upfal, 2017)
@@
Let $Y_{i}^{(t)}\sim\mathrm{Poisson}(tp_{i})$, $i=1,\dots,n$, be independent random variables with $\sum_{i=1}^{n}p_{i}=1$. Moreover, let $(Z_{1}^{(t)},Z_{2}^{(t)},\dots,Z_{n}^{(t)})\sim\allowbreak\mathrm{Multinomial}(t,(p_{1},\dots,p_{n}))$. Let $f:\mathbb{R}^{p}\to\mathbb{R}_{+}$ be any non-negative function. Then
@@
The following propositions are direct consequences of a truncated matrix Bernstein inequality (see e.g., Proposition A.7 in (Hopkins et al., 2016)).
@@
Let $A$ be a $m\times 2r$ deterministic matrix, and $B$ be a $n\times 2r$ deterministic matrix. Then, under the random matrix model (14) with compound Poisson entries, and denoting $L=\|M\|_{\max}\vee\sigma$, we have:
@@
    for all $\ell\in[m]$, for all $\delta\in(0,1)$, for all $T\geq(1/(n\omega_{\min}))\log^{3}(en/\delta)$, the event
@@
    for all $k\in[n]$, for all $\delta\in(0,1)$, for all $T\geq(1/(m\omega_{\min}))\log^{3}(em/\delta)$, the event
@@
The proofs of Proposition B.5 and Proposition B.6 are essentially the same as those of Proposition 26 and Proposition 27 in (Stojanovic et al., 2024), and are omitted. Note that both Propositions in (Stojanovic et al., 2024) were derived under an i.i.d. sub-gaussian noise assumption.
@@
This part of the analysis follows the steps of the proof of Lemma 30 in (Stojanovic et al., 2024), but we repeat the main results for the sake of completeness.
@@
We can define $\widetilde{S}$ in a similar way, and denote $\widehat{Q}\in\mathbb{R}^{(m+n)\times 2r}$, the matrix containing the $2r$ eigenvectors of the best $2r$-rank approximation of $\widetilde{S}$. Here, we observe that:
@@
(ii) Second, it can be shown, under the event
$\mathcal{E}_{1}=\{\|E\|_{\textup{op}}\leq C_{1}\sigma_{r}(M)\}$
where $C_{1}$ is a universal and sufficiently small positive constant, that:
$\|Q-\widehat{Q}(\widehat{Q}^{\top}Q)\|_{2\to\infty}\leq\frac{1}{\sigma_{2r}(S)}\bigg{(}\frac{5\|SQ\|_{2\to\infty}\|E\|_{\textup{op}}}{\sigma_{2r}(S)}+3\|EQ\|_{2\to\infty}+2\|E(Q-\widehat{Q}(\widehat{Q}^{\top}Q)\|_{2\to\infty}\bigg{)}.$ (18)
@@
Now, the key observation that underpins the leave-one-out analysis is that $E_{\ell,:}$ and $Q-\widehat{Q}^{(\ell)}W_{\widehat{Q}^{(\ell)}}$ are independent provided that the entries of the error matrix $E$ are independent. This will be ensured thanks to the Poisson approximation (using Lemma B.3) by controlling the quantities of interest under the random matrix model (14). Define
$\mathcal{B}(\delta)=\sqrt{\frac{m+n}{T\omega_{\min}}\log^{3}\left(\frac{m+n}{\delta}\right)}$
@@
$\mathcal{E}_{2}$ =$=\left\{\|E\|_{\textup{op}}\leq C_{2}L\mathcal{B}(\delta)\right\}$
$\forall\ell\in[m+n],\qquad\mathcal{E}_{3}^{(\ell)}$ =$=\left\{\|E_{\ell,:}(Q-\widehat{Q}^{(\ell)}W_{\widehat{Q}^{(\ell)}})\|_{2}\leq C_{3}L\|Q-\widehat{Q}^{(\ell)}W_{\widehat{Q}^{(\ell)}}\|_{2\to\infty}\mathcal{B}(\delta)\right\}$
@@
for $c>0$ a sufficiently large universal constant. Additionally, using the fact that $L/\sigma_{r}(M)\geq 1/\sqrt{mn}$ from Lemma B.8, we can deduce that $\mathcal{E}_{1}\supseteq\mathcal{E}_{2}$ and $\mathbb{P}(\mathcal{E}_{2})\geq 1-\delta$ as well as $\mathbb{P}(\mathcal{E}_{3}^{(\ell)})\geq 1-\delta$ for all $\ell\in[m+n]$ when
$T\geq\frac{c^{\prime}L^{2}(m+n)}{\sigma_{r}^{2}(M)\omega_{\min}}\log^{3}\left(\frac{m+n}{\delta}\right)$ (21)
for a positive universal constant $c^{\prime}$ sufficiently large. At this point, what remains to be done is to deal with the inequalities (18), (19) and (20) under the intersection of the events $\mathcal{E}_{2},\mathcal{E}_{3}^{(1)},\dots,\mathcal{E}_{3}^{(m+n)}$. We spare the readers the details of this tedious task and write directly the resulting bound
$\|Q-\widehat{Q}(\widehat{Q}^{\top}Q)\|_{2\to\infty}\leq C^{\prime}\frac{L}{\sigma_{r}(M)}\frac{\mu\kappa r^{1/2}\mathcal{B}(\delta)}{\sqrt{m\wedge n}}$
for some universal constant $C^{\prime}>0$. Thus, by a union bound, the above bound holds with probability at least $1-(1+m+n)\delta$, provided that condition (21) holds, under the random matrix model (14). Using Lemma B.3, we conclude that the same event holds with probability $1-(1+m+n)e\sqrt{T}\delta$, provided that condition (21) holds, under the random matrix model (13). Re-parametrizing by $\delta^{\prime}=e(m+n+1)\sqrt{T}\delta$ concludes the proof.
@@ B.4 Subspace recovery in Frobenius norm
Deriving such guarantees is by now well understood under general noise assumptions (see e.g., (Keshavan et al., 2009; Chen et al., 2021)). Indeed, the classical proof follows by first deriving upper bounds on $\|\widetilde{M}-M\|_{\textup{op}}$ with high probability, then using Davis-Kahan’s or Wedin’s theorem to obtain upper bounds on the subspace recovery error in the Frobenius norm. The following lemma provides such upper bounds:
@@
Let $\delta\in(0,1)$ and consider the setting of Theorem 4.1. Then, with probability at least $1-\delta$, it holds that
$\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{\textup{F}},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{\textup{F}}\right)\lesssim\frac{L\sqrt{r}}{\sigma_{r}(M)}\sqrt{\frac{m+n}{T\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)},$
provided that $T\gtrsim\frac{1}{\omega_{\min}(m\wedge n)}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$.
@@
An analogous bound holds for $\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{F}$. Lastly, using Proposition B.5 to bound $\|M-\widetilde{M}\|_{\textup{op}}$ and Lemma B.8 to bound $\|M\|_{\max}/\sigma_{r}(M)$ concludes the proof. ∎
@@ B.5 Additional Lemmas
The following result is immediate (see Lemma 17 in (Stojanovic et al., 2024)).
@@ Appendix C Policy Evaluation for Contextual Linear Bandits
In this appendix, we study the PE task for the simpler contextual linear bandit problem, whose error bound analysis will be useful to handle the PE task for the contextual low-rank bandit problem. We also demonstrate that the linear bandit error bound is instance-optimal by deriving a sample complexity lower bound of independent interest.
We observe $T$ i.i.d. samples of the form $(i_{t},j_{t},r_{t})_{1\leq t\leq T}$ where the context $i_{t}\in[m]$ is drawn independently from a known distribution $\rho$ and the arm $j_{t}\in[n]$ is sampled from $\pi^{b}(.|i_{t})$ for a known behavioral policy $\pi^{b}$. The reward distribution is assumed to have a linear structure with respect to a known feature map $\psi:[m]\times[n]\to\mathbb{R}^{d}$ : there exists an unknown parameter $\theta$ in $\mathbb{R}^{d}$ such that $r_{t}=\psi_{i_{t},j_{t}}^{\top}\theta+\xi_{t}$ where $\left(\xi_{t}\right)_{t\leq 1}$ is an i.i.d. sequence of $\sigma$-subgaussian noise, independent from all the other random variables. The overall bandit model described above will be denoted as $\mathcal{M}_{\theta}$. For a given target policy $\pi$, we use the previously defined notations $\omega_{i,j}=\rho(x)\pi^{b}(j|i)$, $\omega^{\pi}_{i,j}=\rho(x)\pi(j|i)$ and we define
@@
For any $\delta\in(0,1)$ and any $\varepsilon>0$, we say that $\hat{v}^{\pi}$ is $(\varepsilon,\delta)$-PAC [^10] We will often omit the superscript $\pi$ to ease the notations. if $\mathbb{P}_{\theta}\left(|v^{\pi}_{\theta}-\hat{v}^{\pi}|\leq\varepsilon\right)\geq 1-\delta$ for every $\theta\in\mathbb{R}^{d}$. We will write $v^{\pi}$ for $v^{\pi}_{\theta}$ when there is no ambiguity about the underlying parameter. We also write $N\succeq M$ (resp. $N\preceq M$) when $N-M$ (resp. $M-N$) is positive semi-definite, and denote the Moore-Penrose pseudoinverse (Barata & Hussein, 2011) of $N$ by $N^{\dagger}$.